Points A, B, C, D, and E are coplanar and no three are collinear. In how many way can the plane be named using only these points?
Answer by richard1234(7193)

Remember that three or more coplanar points uniquely determine a plane. We can choose three points, or even four or all five points. The number of ways to do so is 5C3 + 5C4 + 5C5 = 10 + 5 + 1 = 16 ways.